OpenStudy (anonymous):

If y varies jointly as x and z, and y = 20, x = 2, and z = 5, then find k, the constant of variation.

OpenStudy (ganpat):

is it directly or inversely ??

OpenStudy (anonymous):

i think it's directly

OpenStudy (ganpat):

so \[y \alpha (x) (z) = y = K (x) (z)\] put all the values.. 20 = k (2) (5) K = 2...